# From Trace to Line: LLM Agent for Real-World OSS Vulnerability Localization

### v2

@@ 4.2 T2L Features
Divergence Tracing. Recognizing that complex vulnerabilities often involves multiple files and functions, T2L-Agent also uses divergence tracing to explore multiple hypotheses in parallel from the same crash log. This feature is inspired by how modern LLM interfaces offer multiple response variations for users, using the LLM’s variability to ensure comprehensive coverage of potential vulnerability locations. Rather than committing to a single chain of thought, it expands several in parallel and returns a ranked list of candidate sites across the search space. This surfaces correct localizations that were not top ranked initially and is especially helpful for bugs that span multiple modules. From Figure 4 (b), we can observe that the divergence tracing generates more localization candidates and matched more vulnerable lines in this round.
Detection Refinement. The detection refinement begins based on crash logs and initial localization candidates. Each LLM-generated candidate gets a confidence label from very unlikely to very likely. We use these labels to sort and rank the candidates during the coarse-grained filtering. T2L agent selects code slices based on crash logs, stack trace information, and vulnerability patterns instead of exhaustively examining the full code snippet. In fine-grained refinement, we extract these slices and feed them back to the LLM along with the unified evidence, providing richer context surrounding the initially flagged locations.
The refinement process operates iteratively to help the agent correct early mistakes and discover vulnerabilities that are not immediately obvious from traces, particularly for complex vulnerabilities involving memory corruption where the crash point may be far away from the actual vulnerability. As Figure 4 (a) shows, based on the source code LLM interested, the refinement process successfully locates new lines that were not found during first step with only the runtime evidence.
@@ 4.3 T2L-ARVO Benchmark
The final benchmark comprises 50 vulnerabilities samples, evenly sampled across five crash families (10 each) for broad yet controlled difficulty. Each family includes representative subtypes (e.g., heap-buffer-overflow, use-of-uninitialized-value, heap-use-after-free, bad-cast), covering single-file defects and cross-module interactions to prevent overflow bias and exercise diverse failure modes observed in real repositories.
Figure 5: Crash types in T2L-ARVO Bench.
@@
Verification Process Our verification follows two main considerations. First, ARVO crash families contain multiple subtypes. To reduce redundancy and improve representativeness, we select only a few typical subtypes from each family (see A.1), preserving diversity without overloading the benchmark with similar vulnerabilities. Second, experts evaluate each sample’s complexity to keep T2L-ARVO at a moderate, calibrated difficulty. We score candidates using diff-based structural metrics (e.g., files changed, architectural spread, directory depth) and semantic factors (e.g., cross-module coupling, interface changes), ensuring that T2L-ARVO remains balanced in both coverage and difficulty.
@@ 5 Experiment Setup
Metrics. We report two complementary scores for project-level OSS vulnerability studies. Detection asks whether the agent flags a vulnerability within the correct module/chunk and materially shrinks the search space. Localization requires exact line matches to ground-truth patches. Together, they separate “finding the neighborhood” from “pinpointing the line”, mirroring real debugging.
Data Preparation. We evaluate on the full T2L-ARVO set: 50 verified, structured challenges derived from ARVO, spanning diverse domains (e.g., imaging, networking) and balanced complexity so results reflect production patterns rather than a single project or bug family. Because ARVO lacks detection-ready chunking, T2L-ARVO adds AST-based segmentation: projects are partitioned into semantically meaningful units for scoring coarse detection, while exact line matches assess fine localization—a single framework for both levels.
Model Selection. We assess a set of state-of-the-art language models—both open-source and commercial to probe generality and robustness of T2L-Agent across architectures and scales, including open models such as Qwen3 Next, Qwen3 235B, DeepSeek 3.1, LLaMA 4 and commercial models like Claude4 Sonnet, GPT-5, GPT-4.1, GPT-4o-mini, Gemini 2.5 Pro, Gemini 2.5 Flash with a maximum budget \$1.0. We use API keys from commercial model’s official providers and Together.ai’s inference service for open source models.
Implementation. We build T2L-Agent from scratch without LangChain, DSPy, and LlamaIndex to keep the core lightweight, retain fine-grained control over reasoning and tools, and maximize extensibility. The Agentic Trace Analyzer compiles targets with ASAN and collects crashes, stack traces, and allocation metadata to yield actionable traces for narrowing. Following ARVO’s layout, we maintain registry-backed, per-project environments and provide both vulnerable and patched revisions in containers. Our harness runs dockerized T2L-Agent’s that interface with T2L-ARVO via the Docker SDK for Python, orchestrating build–run–reproduce cycles from within the agent loop, ensuring consistent conditions, deterministic reproduction, and auditable measurement.
@@ 6.1 Baseline Benchmarking
We evaluate T2L-Agent on T2L-ARVO end-to-end and report Detection Rate and Localization Rate. Table 1 covers five models—GPT-5, GPT-4.1, GPT-4o-mini, Qwen 3 Next 80B, and Qwen3 235B, running under identical per-case budgets, environments, and AST-based chunking. Overall, detection is higher than localization by design. Under this setting, GPT-5 leads localization at 41.7%, followed by GPT-4.1 at 38.5%; GPT-4o-mini lands at 22.6%, and open-source models trail Qwen 3 235B 9.2% and Qwen 3 Next 80B 5.9%. For detection, GPT-4.1 is highest at 48.0%, with GPT-5 and GPT-4o-mini both at 44.3%; Qwen3 reaches 37.4% and Qwen3 235B 25.9%. While Gemini2.5 Pro shows limited effectiveness with 17.4% detection and 10.5% localization rates.
Family-wise patterns are consistent across metrics. Buffer and Memory are easier due to concrete runtime cues: for localization, GPT-5 reaches 53.8% and 55.9%, and most models cluster near the mid-50s for detection. Initialize sits mid-range and benefits from multi-step reasoning (e.g., GPT-5 35.5% vs. GPT-4.1 46.4% in localization). Parameter is often solvable from interface/call-site context (39.4% for GPT-5; 29.9% for GPT-4.1). Runtime remains uniformly hardest: detection hovers around 11.1–21.2% even for top configurations, and the best localization we observe is 20.5% on GPT-4.1, reflecting sparse, unstable traces.
Taken together, equal budgets surface clear profiles. GPT-5 (low-think) converts evidence into the strongest line-level localization, while GPT-4.1 extracts slightly more coarse-grained signal at the chunk level. GPT-4o-mini’s mix—competitive detection (44.3%) but weak localization (22.6%), suggests higher recall with looser ranking that does not always translate to precise line hits. Open-source models lag on both metrics under the same constraints, indicating gaps in code understanding and tool use rather than simple parameter tuning. Overall, improvements track the availability of concrete runtime evidence, and structured, tool-grounded reasoning appears more impactful than generation settings for end-to-end vulnerability localization.
@@ 6.2 Discussion 1: Feature-wise Evaluation
Agentic Trace Analyzer. This table  2 demonstrates the critical effectiveness of our proposed Agentic Trace Analyzer (ATA) through ablation experiments. Without ATA, GPT-5 and Claude 4 Sonnet show 0.0% detection and localization across all families. This performance breakdown validates that our ATA component successfully bridges the gap between crash symptoms and vulnerability locations, addresses the fundamental challenge of vulnerability localization in complex codebases.
Detection Refinement. Compared with Tab  1, Tab. 2 shows broad, across-the-board gains after enabling refinement. Strong proprietary models improve steadily, while open-source models jump the most—Qwen3 235B’s localization rises by roughly sevenfold. Improvements vary by crash family: Initialize bugs benefit most (they demand multi-step reasoning), whereas Buffer and Memory see smaller lifts because concrete runtime evidence already anchors the search. Runtime cases remain hard—when traces are sparse, refinement offers limited benefit. Net effect: higher recall and more precise line-level hits with minimal tuning. Several additional models show promising performance. Deepseek V3.1 achieves the highest overall results with 53.9% detection and 53.4% localization rate. LLaMa 4 demonstrates balanced capabilities on both metrics, and Gemini 2.5 Flash shows variable performance across crash families.
@@ 6.3 Discussion 2: Parameter Tuning
Thinking Budget. As shown in Tab.  3, more thinking didn’t help. On GPT-5, the Medium budget outperforms High with 50.9% detection vs 41.3%, and 41.6% localization vs 36.1%. The Low setting trails Medium by a few points yet often matches or even exceeds High on key metrics, while sharply reducing compute and latency. This pattern suggests diminishing returns—and decision drag—at very high budgets: the model over-explores, delays commitment, and accumulates tool-use errors. In practice, Medium strikes the best accuracy–cost balance for vulnerability localization; Low is a strong option when throughput and responsiveness matter most.
@@
Temperature. Temperature changes barely matter from Tab.  3. On GPT-4.1, detection is 51.0% at 0.2 and 50.8% at 0.6, with localization 43.6% vs. 43.5%. Claude 4 Sonnet shows the same pattern: 46.5% vs. 47.3% detection and 43.0% vs. 44.9% localization. Performance is stable in 0.2-0.6 range. The exception is Initialize bugs, which are more temperature sensitive than Buffer and Memory cases that lean on concrete runtime evidence. Overall, precise localization benefits more from structured, tool-grounded reasoning than from extra sampling, making parameter choices simple.
6.4 Discussion 3: Case Study
@@ A.3 Model Failure Analysis
We conducted a targeted failure analysis on several baseline models to map out common failure modes as shown in Tab.  12. Claude 4 Sonnet and Gemini 2.5 Pro hit the budget ceiling in 81.6% and 85.7% of runs respectively, indicating efficient resource utilization. GPT-5 reaches 61.2% with execution errors (28.6%), while Qwen3 235B struggles with basic data operations (59.2%). Open-source baselines stall early: Qwen 3 Next fails to surface actionable candidates in 44.9% of trials. Execution errors remain common across older models (20–30%), showing that tool use often breaks even when a plan exists. Net-net, while newer models show improved resource management, legacy models skew toward either incomplete exploration or difficulty navigating real-world code.

### v3

@@ 4.2 T2L Features
Divergence Tracing. Recognizing that complex vulnerabilities often involves multiple files and functions, T2L-Agent uses divergence tracing to improve coverage. Rather than committing to a single chain of thought, it explores multiple parallel reasoning branches from the same evidence block and aggregates them into a ranked list of candidates This helps recover correct localizations that may not appear in the top hypothesis of a single run and is especially helpful for bugs that span multiple modules. From Figure 2 (a), we can observe that the divergence tracing generates more localization candidates and matched more vulnerable lines in this round.
Detection Refinement. To reflect the coarse-to-fine nature of the localization task, T2L baseline solver uses a iterative two-stage refinement loop. It begins with coarse-grained filtering: the system derives broad candidate regions from crash logs, using LLM code comprehension to link symptoms to likely causes and produce a ranked list of suspicious location candidates with confidence and rationale. Each LLM-generated candidate gets a confidence label from very unlikely to very likely. We use these labels to sort and prioritize candidates for subsequent inspection.
Instead of exhaustively scanning long code contexts, T2L selects targeted code slices guided by the crash log, stack trace information, and common vulnerability patterns such as missing bounds checks, uninitialized variables, or improper memory management. Concretely, we extract code fragments that the LLM explicitly flagged as relevant in the first pass, and enter the fine-grained refinement stage: these extracted fragments are appended back to the unified evidence as additional context providing richer context surrounding the initially flagged locations for a second LLM pass, which updates the candidate set. Newly proposed locations are merged with existing candidates. This refinement loop continues until no new candidates emerge or the cost budget is reached.
The refinement process operates iteratively to help the agent correct early mistakes and discover vulnerabilities that are not obvious from traces, particularly for complex vulnerabilities involving memory corruption bugs where the crash point can be far from the actual vulnerability. As Figure 2(b) shows, incorporating the LLM-identified source snippets enables refinement to discover new vulnerable lines that were missed in the initial pass using only the runtime evidence.
@@ 4.3 T2L-ARVO Benchmark
The final benchmark comprises 50 vulnerabilities samples, evenly sampled across five crash families (10 each) for broad yet controlled difficulty. Each family includes representative subtypes, covering single-file defects and cross-module interactions to prevent overflow bias and exercise diverse failure modes observed in real repositories.
Table 2: Crash types in T2L-ARVO Bench.
@@
Verification Process Our verification follows two main considerations. First, ARVO crash families contain multiple subtypes. To reduce redundancy and improve representativeness, we select only a few typical subtypes from each family (see A.1), preserving diversity without overloading the benchmark with similar vulnerabilities. Second, experts evaluate each sample’s complexity to keep T2L-ARVO at a moderate, calibrated difficulty. We score candidates using diff-based metrics (e.g., files changed, architectural spread, directory depth) and semantic factors (e.g., cross-module coupling, interface changes), ensuring that T2L-ARVO remains balanced in both coverage and difficulty.
@@ 5 Experiment Setup
Metrics. We report two complementary scores for project-level OSS vulnerability studies. Detection asks whether the agent flags a vulnerability within the correct module/chunk. Localization requires exact line matches to ground-truth patches. Together, they separate ”finding the neighborhood” from ”pinpointing the line”, mirroring real debugging.
Data Preparation. We evaluate on the full T2L-ARVO set: 50 verified challenges derived from ARVO, spanning diverse domains (e.g., imaging, networking) and balanced complexity so results reflect production patterns. Because ARVO lacks detection-ready chunking, T2L-ARVO adds AST-based segmentation: projects are partitioned into semantically meaningful units for scoring coarse detection, while exact line matches assess fine localization.
Model Selection. We assess state-of-the-art language models, both open-source and commercial, to probe generality and robustness across architectures and scales, including Qwen3 Next, Qwen3 235B, DeepSeek 3.1, LLaMA 4, Claude4 Sonnet, GPT-5, GPT-4.1, GPT-4o-mini, Gemini 2.5 Pro, and Gemini 2.5 Flash with a maximum budget \$1.0. We use API keys from official providers and Together.ai for open-source models.
Implementation. We build T2L-Agent from scratch without LangChain, DSPy, or LlamaIndex to keep the core lightweight and retain fine-grained control. The Agentic Trace Analyzer compiles targets with ASAN and collects crashes, stack traces, and allocation metadata. Following ARVO’s layout, we maintain per-project environments and provide both vulnerable and patched revisions in containers. Our harness runs dockerized T2L-Agent’s via the Docker SDK for Python, orchestrating build–run–reproduce cycles and ensuring deterministic reproduction.
Comparison with existing localization agent architectures. We compare against representative agent architectures from AgentFL and AutoFL. Rather than reusing their original systems, we implement architecture preserving transfer baselines with minimal adaption mappings to our crash-driven trace-to-line protocol, including replacing test signals with trace-derived evidence and generating line predictions for unified scoring.
@@ 6.1 Baseline Benchmarking
We evaluate T2L-Agent on T2L-ARVO and report Detection Rate and Localization Rate. Table 3 covers seven models: GPT-5, GPT-4.1, GPT-4o-mini, Claude 4 Sonnet, Gemini2.5 Pro, Qwen3 235B, and Qwen 3 Next 80B, running under identical per-case budgets, environments, and AST-based chunking. Overall, detection is higher than localization. Under this setting, GPT-5 leads localization at 41.7%, followed by GPT-4.1 at 38.5% and Claude 4 Sonnet at 30.5%; open-source models trail below 10%. For detection, GPT-4.1 is highest at 48.0%, with GPT-5 and Claude 4 Sonnet close behind; Gemini2.5 Pro shows limited effectiveness at 17.4%.
Family-wise patterns are consistent. Buffer and Memory are easier due to concrete runtime cues: GPT-5 reaches 53.8% and 55.9% localization, and most models cluster near the mid-50s for detection. Initialize sits mid-range and benefits from multi-step reasoning. Parameter is often solvable from interface/call-site context. Runtime remains uniformly hardest: detection hovers around 11–21% even for top configurations, reflecting sparse, unstable traces.
Taken together, equal budgets surface clear profiles. GPT-5 converts evidence into the strongest line-level localization, while GPT-4.1 extracts slightly more coarse-grained signal at the chunk level. GPT-4o-mini shows competitive detection but weak localization, suggesting higher recall that does not always translate to precise line hits. Open-source models lag on both metrics, indicating gaps in code understanding and tool use. Overall, improvements track the availability of concrete runtime evidence, and structured, tool-grounded reasoning appears more impactful than generation settings for end-to-end vulnerability localization.
@@ 6.2 Discussion 1: Feature-wise Evaluation
Agentic Trace Analyzer. This table  4 demonstrates the critical effectiveness of our proposed Agentic Trace Analyzer (ATA) through ablation experiments. Without ATA, GPT-5 and Claude 4 Sonnet show 0.0% detection and localization across all families. This performance breakdown validates that our ATA component successfully bridges the gap between crash symptoms and vulnerability locations, addresses the fundamental challenge of vulnerability localization in complex codebases.
Like human debugging, T2L-Agent baseline solver seeds candidates via static–dynamic correlation and iteratively refines them with source–feedback loops. ATA brings fewer single shot failures, improved compute efficiency, and behavior based decisions that enable precise, line-level localization even in large, tightly coupled codebases. With ATA disabled, the LLM localizes on its own, as shown in Figure 2(c).
Detection Refinement. Compared with Tab  3, Tab. 4 shows broad, across-the-board gains after enabling refinement. Strong proprietary models improve steadily, while open-source models jump the most—Qwen3 235B’s localization rises by roughly sevenfold. Improvements vary by crash family: Initialize bugs benefit most (they demand multi-step reasoning), whereas Buffer and Memory see smaller lifts because concrete runtime evidence already anchors the search. Runtime cases remain hard—when traces are sparse, refinement offers limited benefit. Net effect: higher recall and more precise line-level hits with minimal tuning. Several additional models show promising performance. Deepseek V3.1 achieves the highest overall results with 53.9% detection and 53.4% localization rate. LLaMa 4 demonstrates balanced capabilities on both metrics, and Gemini 2.5 Flash shows variable performance across crash families.
@@ 6.3 Discussion 2: Parameter Tuning
Thinking Budget. As shown in Tab.  5, more thinking didn’t help. On GPT-5, the Medium budget outperforms High with 50.9% detection vs 41.3%, and 41.6% localization vs 36.1%. The Low setting trails Medium by a few points yet often matches or even exceeds High on key metrics, while sharply reducing compute and latency. This pattern suggests diminishing returns—and decision drag—at very high budgets: the model over-explores, delays commitment, and accumulates tool-use errors. In practice, Medium strikes the best accuracy–cost balance for vulnerability localization; Low is a strong option when throughput and responsiveness matter most.
@@
Temperature. Temperature changes barely matter from Tab.  5. On GPT-4.1, detection is 51.0% at 0.2 and 50.8% at 0.6, with localization 43.6% vs. 43.5%. Claude 4 Sonnet shows the same pattern: 46.5% vs. 47.3% detection and 43.0% vs. 44.9% localization. Performance is stable in 0.2-0.6 range. The exception is Initialize bugs, which are more temperature sensitive than Buffer and Memory cases that lean on concrete runtime evidence. Overall, precise localization benefits more from structured, tool-grounded reasoning than from extra sampling, making parameter choices simple.
Figure 3: T2L-Agent pipeline visualization.
6.4 Discussion 3: Compared Frameworks
We select AgentFL and AutoFL as representative agent-based localization frameworks combining LLM-guided navigation with code comprehension, making them natural baselines for our repo-scale trace-to-line task.
For AgentFL, we follow its three-stage workflow (fault comprehension, codebase navigation, fault confirmation) with minimal changes: replacing test-based inputs with crash logs, guiding navigation over relevant files, and applying fault confirmation via view source tools. The baseline outputs the top-1 method plus a line span. For AutoFL, we follow its two-stage design. In Stage 1, the agent queries covered classes (source files) and methods (functions), retrieving code for selected candidates. In Stage 2, navigation tools are locked and the agent outputs suspicious methods with line spans. Both baselines use identical evaluation metrics.
Table 6 shows T2L consistently outperforms both transferred frameworks across models in detection and localization rate, indicating that current agent pipelines struggle to turn coarse detection into reliable line-level localization.
AgentFL achieves acceptable detection but substantially lower localization than T2L. Its best results come from GPT-5 (29.8%) and GPT-4.1 (26.2%), compared with 41.7% and 38.5% for T2L. However, trends vary: for Qwen models, AgentFL reaches slightly higher localization than T2L, suggesting existing pipelines sometimes work but behave inconsistently across models. AutoFL’s performance drops sharply—localization remains low (27.0%) and near zero for smaller models like Qwen3 Next 80B (2.6%).
These results suggest that while coarse chunks can sometimes be detected, accurately narrowing to vulnerable lines remains difficult without tailored trace-to-line methods. Effective systems must use runtime evidence to connect symptoms to root-cause lines. Overall, transferred baselines reach substantially lower performance than T2L, indicating current agent architectures are insufficient for crash-driven, line-level localization in complex repositories.
T2L AgentFL AutoFL
Model Det Loc Det Loc Det Loc
GPT-5 44.3 41.7 33.1 29.8 33.9 27.0
GPT-4.1 48.0 38.5 27.3 26.2 18.3 9.0
GPT-4o-mini 44.3 22.6 21.6 13.8 6.0 4.0
Claude 4 Sonnet 45.9 30.5 38.8 25.4 29.8 13.3
Gemini2.5 Pro 17.4 10.5 17.1 10.3 12.3 3.5
Qwen3 235B 25.9 9.2 14.6 12.5 20.3 8.7
Qwen3 Next 80B 37.4 5.9 11.0 7.5 11.5 2.6
Table 6: Average detection (Det) and line-level localization (Loc) rates on T2L-ARVO (%).
6.5 Discussion 4: Case Study
@@ A.3 Model Failure Analysis
We conducted a targeted failure analysis on several baseline models to map out common failure modes as shown in Tab.  9. Claude 4 Sonnet and Gemini 2.5 Pro hit the budget ceiling in 81.6% and 85.7% of runs respectively, indicating efficient resource utilization. GPT-5 reaches 61.2% with execution errors (28.6%), while Qwen3 235B struggles with basic data operations (59.2%). Open-source baselines stall early: Qwen 3 Next fails to surface actionable candidates in 44.9% of trials. Execution errors remain common across older models (20–30%), showing that tool use often breaks even when a plan exists. Net-net, while newer models show improved resource management, legacy models skew toward either incomplete exploration or difficulty navigating real-world code.